Find the density of a hydrogen nucleus in grams per cubic centimeters,

given that the diameter of this nucleus is 1.0*10 pm and the mass is 1.67

* 10-24g. (Hint Volume of a sphere of radius r is far.)


1
Expert's answer
2021-05-10T01:50:30-0400

"Diameter = 10 pm = 1\u00d7 10^{-11}M = 1\u00d710^{-9}cm"

"Mass= 1.67\u00d710^{-24}g"

"V= \frac{4}{3}× 3.14 × [\frac{1}{2}×(1×10^{-9})]^3 = 4.2×10^{-8}cm3"

"D= \frac{1.67×10^{-24}}{4.2×10^{-8}} = 4.04× 10^{-17}g/cm³"